# Advances in Kidney Biopsy Lesion Assessment through Dense Instance Segmentation

### v2

@@ 3.1 Preliminaries
Diffusion model. Diffusion approaches ([43, 39]) is a family of deep generative models, inspired by the principles of non-equilibrium thermodynamics ([42, 25]). Their operation is conceptualized as $T$-step sequential process, iteratively transiting from an initial state $Z_{0}$ to a final state of pure noise $Z_{T}$. Diffusion forward involves the progressive addition of Gaussian noise at each transition. Due to the special properties of the Gaussian distribution ([24]), it is possible to directly sample any intermediate noisy at state $Z_{t}$:
@@
where $Z_{0}$ is the original data, $Z_{t}\ (t\leq T)$ is the latent state at step $t$, and $\bar{\alpha}_{t}$ is the cumulative noise variance schedule ([8]). Notably, the diffusion forward process introduces no trainable parameters. The objective of a diffusion model is to learn the reverse process. A neural network, $f_{\theta}(Z_{t},t)$, approximates the recovery of the original distribution $Z_{0}$ from a noisy state $Z_{t}$. Training optimizes this network by minimizing a loss function to denoise at arbitrary intermediate state $t$:
@@
where $Z_{0}$ represents the ground-truth and $Z_{t}$ is the noisy input. During inference, the model $f_{\theta}$ reconstructs from pure noise $Z_{T}$ to the predicted data $Z_{0}$ through a series of reverse updates ([43]) with a step-size $s$: $Z_{T}\rightarrow Z_{T-s}\rightarrow\cdots\rightarrow Z_{0}$. A detailed derivation is provided in ([43]).
@@ 3.2 Dense Instance Segmentation Model
Due to the iterative proposal generation, DiffRegFormer necessitates multiple model executions at inference, posing a computational burden if the entire model runs on the raw image at each step during training. To address this issue, we decouple the image encoder from the remaining modules. This strategy enables the encoder to extract multi-scale feature maps from the raw input image only once. Then subsequent modules operate for iterative refinement of box and mask predictions from an initial state of Gaussian noise conditioned on these deep features. Moreover, to facilitate the mask-decoder training, we introduce a novel sampling module that randomly selects positive samples from each class, ensuring a balanced distribution of class-wise regional features. This, thereby, facilitates the learning of object representations with lower frequencies in the dataset.
@@
    Encoder: Processes the raw input image once to extract multi-scale feature maps for bbox-decoder and mask-decoder seperately.
@@
    Sampling Module: Stablizes the training by randomly selecting class-wise balanced positive samples for the mask-decoder .
@@ 3.2.1 Encoder
The image encoder extracts high-level feature maps from the raw image for subsequent decoders. It utilizes a ResNet pre-trained on ImageNet ([22]) as its backbone, followed by a Feature Pyramid Network (FPN) ([32]) to generate multi-scale feature maps. After FPN, a stack of $3\times 3$ convolutional operations is applied to produce two separate sets of feature maps, thereby decoupling the bbox-decoder and mask-decoder.
3.2.2 Bbox decoder
The bbox-decoder builds upon DiffusionDet ([8]). As shown in Figure.4(b), it firstly takes a set of random boxes and multi-scaled feature maps from the encoder as inputs. Dynamic queries (Figure.4(a)) then are initialized from a regional pooling operator (RoIAlign ([21])) and a feed-forward network (FFN) ([46]). These queries undergo iterative refinement through multiple stages. Each stage takes proposal boxes and dynamic queries from the previous stage, generating refined dynamic queries, box predictions, and class predictions for the next stage. Figure.4(c) illustrates one refinement stage in details. Firstly new regional features (cropped by RoIAlign) interact with dynamic queries (after a self-attention module) via dynamic convolution ([11]), emphasizing regions likely containing objects while suppressing others. Hereafter, the enhanced regional features are processed by FFNs to produce refined boxes, box-wise classifications, and refined queries respectively.
@@ 3.2.3 Sampling
Conventional RCNN models use intersection over union (IoU) to classify a proposal box as positive if its IoU with a ground-truth box is $\geq 0.5$ ([21]). This approach, however, is ill-suited for diffusion methods, particularly in early training stages when diffusion methods produce predominantly noisy. This can hinder convergence or even cause failure due to error accumulation. Further, conventional IoU-based methods are susceptible to class imbalances, biasing the learning process towards frequently occurring objects, thereby neglecting rare instances. We propose a novel sampling method to address these issues. We firstly replace proposal boxes with ground-truth boxes, ensuring a sufficient supply of positive samples. This is justified because the bbox-decoder’s objective is to predict boxes that closely approximate ground truth. We then divide ground-truth bounding boxes into groups based on their classes. Within each group, we randomly select up to $n$ boxes as positive samples without repetition. This sampling strategy offers three distinct advantages:
@@
    Randomize Frequent Instance: Learning on a randomized subset of frequent instances prevents model overfitting.
@@ 3.2.4 Mask decoder
The mask decoder refines mask predictions iteratively, utilizing regional features cropped from multi-scale feature maps based on boxes selected by the sampling module. Instead of using global queries ([12, 13]), we compute cross-attention between dynamic queries and positive regional features. This highlights effective queries and filters others, enhancing regional features by focusing on pixels within positive regions, thereby facilitating accurate per-pixel instance score prediction (Figure 5).
Fig. 4: The bbox-decoder takes as input multi-scale feature maps and a set of proposal boxes and outputs the prediction of classes and boxes iteratively. (a) The orange rectangle in (b) generates the initial dynamic queries via a RoIAlign pooling operator and a feed-forward network (FFN). (b) The module consists of a dynamic queries initialization head and multiple boxes refinement heads. (c) One box refinement module (one blue rectangle in (b)) inputs the dynamic queries and proposal boxes from the previous stage and generates predictions and refined dynamic queries for the next stage.
Fig. 5: The mask-decoder takes multi-scale feature maps and a set of positive proposal boxes to predict instance masks. The dynamic queries interact with regional features (i.e. cross-attention) and only highlight pixels residing in proposal boxes. Final instance masks are generated from the enhanced regional feature maps.
Fig. 6: The lesion classifier consists of 2 prediction heads for glomeruli and tubuli respectively. It takes crops of anatomical structures and outputs the probability of class-wise lesions.
Fig. 7: An illustration for a diffusion model at different stages. For simplicity, there are only 5 states with skip step as 1. State 4 is pure Gaussian noise while state 0 is the ground truth. (a) In one instance of training process, the diffusion model only randomly selects several denoise paths to train the box decoder; (b) In one instance of inference process, the diffusion model can alternately pick between denoise and corrupt path to generate multiple predictions, and ensemble them together for the final refined results.
@@ 3.3 Lesion classifier
The lesion classifier comprises multiple independent lesion prediction heads. Each head specializes in a specific anatomical structure (e.g., glomeruli) and processes cropped images of the specified structure to predict probabilities for multi-label lesions. Figure.6 illustrates this design. The architecture of a single head consists of three convolutional blocks followed by a max-pooling operator. Each block contains a $3\times 3$ convolutional, batch normalization, and ReLU activation. We currently implement two prediction heads: one for glomeruli (predicting sclerosis) and one for tubuli (predicting atrophy).
Importantly, the modular design of the instance segmentation component allows lesion prediction heads to operate as independent plug-ins. This modularity is crucial for seamlessly adapting the model to the expanding datasets, particularly in clinical scenarios. It allows for the straightforward integration or removal of prediction heads with minimal impact on the overall model architecture. For example, if new annotations for arteries become available, a dedicated prediction head can be trained to classify arterial lesions and integrated into the model without modifying existing components. Likewise, this design allows for the extension of the model’s diagnostic scope to include additional lesions within a given structure (e.g., tubuli) through the simple replacement of an existing prediction head with a newly trained one on expanded data. This design promotes the model’s versatility and long-term adaptability within clinical scenarios where the scope of identifiable lesions may increase.
@@ 3.4 Implementation Details
The entire model adopts distinct strategies for training and inference phases. During training, the DiffRegFormer learns to adeptly reconstruct ground truth boxes at state 00 from an arbitrary noisy state $t$ (Figure.7.(a)). Additionally, the sampling module provides class-balanced ground-truth boxes to facilitate the mask decoder’s training. Lastly, The lesion classifier operates on crops generated from ground-truth instance masks, with each prediction head specializing in classifying lesions specific to the corresponding anatomical structure within these crops. Conversely, during inference, the DiffRegFormer assembles multiple predictions to refine final results (Figure 7(b)). Similarly, The sampling module is deactivated, and all proposals are directly forwarded to the mask decoder for instance segmentation prediction. Likewise, the lesion classifier utilizes instance masks predicted by DiffRegFormer to crop structures. Lastly, each prediction head predicts probabilities for class-specific lesions within these cropped regions.
@@ 3.4.1 Training
We first construct noisy boxes from the ground truth. The DiffRegFormer is then trained to recover bounding boxes and iteratively generate instance masks from an arbitrary state $t$. The algorithm is detailed in Figure.3.2.4.
Diffusion forward. Ground-truth boxes are first concatenated with additional random boxes to aligned to a fixed length across images, accounting for the variability in instance counts. Gaussian noise is then introduced the noise scale governed by $\bar{\alpha}$ as per Eq.(1). Following ([36]), the scale varies with state $t$ using a monotonically decreasing cosine schedule. Notably, for optimal performance, the signal-to-noise ratio requires a relatively high signal scaling value compared to image generation tasks ([9, 24, 16, 10, 8]).
Training Loss. The box decoder takes noisy boxes as input and predicts category classifications and box coordinates. Given the one-to-many relationship between ground-truth and predicted boxes, a set-loss function is firstly employed to match predictions to ground-truths based on the lowest cost ([5, 45, 50]). Furthermore, the mask decoder takes sampled positive boxes as input and predicts instance masks. Due to one-to-one mapping, we then use a Binary Cross Entropy loss function. Finally, since multiple lesions can exist within one crop, a Binary Cross Entropy loss function is also appled to multi-label prediction within each head.
Sampling. Our implementation uses $3$ classes and a maximum of $N=500$ objects per image. we set up to $n=\frac{500}{3}\simeq 133$ instances per class. More specifically, positive ground-truth boxes are sampled without repetition, not exceeding the class-wise maximum. Moreover, we disregard absent classes within the input. For instance, if there are 2 glomeruli, 0 vessels, and 200 tubuli ground-truth boxes within an ROI, the sampling module then would select 2 glomeruli, 0 vessels, and 133 random tubuli as positive samples, without duplication.
@@ 3.4.2 Inference
During inference, the DiffRegFormer starts from a standard Gaussian distribution corresponding to the final state $T$ as defined in Eq.(1). Then a progressive denoising operation reverses the predictions to the initial state 00. Algorithm (as in Figure.3.4.2) outlines this process.
Inference steps. Each inference iteration involves two primary operations. Firstly, the box decoder processes noisy boxes generated from the previous state $\mathrm{t_{now}}$ and predicts both categories and box coordinates. Subsequently, Denoising Diffusion Implicit Models (DDIM) ([35, 43]) introduce noise into the previously predicted boxes. This step generates new noisy boxes for subsequent state $\mathrm{t_{next}}$, facilitating a progressive denoising process across different states. Notably, each iteration performs a single denoising operation, transitioning from state $\mathrm{t_{now}}$ to $\mathrm{t_{next}}$. This characteristic allows for the assembly of intermediate predictions from multiple denoising steps (i.e. $t_{i}\to 0$ where $i\in[1,T]$) refine the final outcome. However, to improve efficiency, DiffRegFormer performs inference only once, directly transitioning from pure Gaussian noise at state $T$ to state 00, to obtain the final predictions.
Box replacement. Predicted boxes can be categorized as either positive (scores above a threshold, containing objects of interest) or negative (scores below the threshold, arbitrarily located). Direct applying negative boxes into DDIM would significantly degrade the quality of newly generated noisy boxes. This issue stems from the fact that negative boxes, originating from corrupted boxes during the training phase, deviate significantly from the Gaussian distribution. To ensure consistency between training and inference phases, we substitute negative boxes with random boxes sampled from a Gaussian distribution.
Table 1: Overall evaluation result
@@
  • 1.
    All models use ResNet-50 as a backbone network, and trained 40000 iterations on one NVIDIA GeForce RTX 3090 GPU with batch size 2. APS, APM, and APL correspond to AP on small, medium, and large-sized objects respectively.
Table 2: Per class evaluation result
Bounding Boxes - Glomeruli Instances - Glomueruli
@@
Cascade Mask-RCNN 23.0 38.0 26.9 - 17.8 33.3 20.1 39.4 21.9 - 16.9 27.5
our model DiffRegFormer 25.7 46.9 27.8 13.5 23.9 31.6 23.4 48.8 20.2 20.2 20.4 30.1
@@
  • 1.
    All models use ResNet-50 as a backbone network, and trained 40000 iterations on one NVIDIA GeForce RTX 3090 GPU with batch size 2. APS, APM, and APL correspond to AP on small, medium, and large-sized objects respectively.

### v3

@@ 3.1 Preliminaries
Diffusion model. Diffusion approaches [53, 49] is a family of deep generative models inspired by the principles of non-equilibrium thermodynamics [52, 30]. Their operation is conceptualized as $T$-step sequential process, iteratively transiting from an initial state $Z_{0}$ to a final state of pure noise $Z_{T}$. Diffusion forward involves the progressive addition of Gaussian noise at each transition. Due to the special properties of the Gaussian distribution [29], it is possible to sample any intermediate noisy at state $Z_{t}$ directly:
@@
where $Z_{0}$ is the original data, $Z_{t}\ (t\leq T)$ is the latent state at step $t$, and $\bar{\alpha}_{t}$ is the cumulative noise variance schedule [8]. Notably, the diffusion forward process does not introduce trainable parameters. The objective of a diffusion model is to learn the reverse process. A neural network, $f_{\theta}(Z_{t},t)$, approximates the recovery of the original distribution $Z_{0}$ from a noisy state $Z_{t}$. Training optimizes this network by minimizing a loss function to denoise at arbitrary intermediate state $t$:
@@
where $Z_{0}$ represents the ground-truth and $Z_{t}$ is the noisy input. During inference, the model $f_{\theta}$ reconstructs from pure noise $Z_{T}$ to the predicted data $Z_{0}$ through a series of reverse updates [53] with a step-size $s$: $Z_{T}\rightarrow Z_{T-s}\rightarrow\cdots\rightarrow Z_{0}$. See [53] for a detailed derivation.
@@ 3.2 Dense Instance Segmentation Model
The iterative proposal generation in DiffRegFormer necessitates multiple model executions at inference. That poses a computational burden if the entire model runs on the raw image at each step during training. To address this issue, we decouple the image encoder from the remaining modules. This strategy enables the encoder to extract multi-scale feature maps from the raw input image only once. Then, subsequent modules operate for iterative refinement of box and mask predictions from an initial state of Gaussian noise conditioned on these deep features. Moreover, to facilitate the mask-decoder training, we introduce a novel sampling module that randomly selects positive samples from each class, ensuring a balanced distribution of class-wise regional features. That facilitates learning object representations with lower frequencies in the dataset.
@@
    Encoder: Processes the raw input image once to extract multi-scale feature maps for bbox-decoder and mask-decoder separately.
@@
    Sampling Module: Stabilizes the training by randomly selecting class-wise balanced positive samples for the mask-decoder.
@@ 3.2.1 Encoder
The image encoder extracts high-level feature maps from the raw image for subsequent decoders. To generate multi-scale feature maps, it utilizes a ResNet [27] pre-trained on ImageNet [15] as its backbone, followed by a Feature Pyramid Network (FPN) [40]. After FPN, a stack of $3\times 3$ convolutional operations is applied to produce two separate feature maps, thereby decoupling the bbox-decoder and mask-decoder.
3.2.2 Bbox-decoder
The bbox-decoder builds upon DiffusionDet [8]. As shown in Figure 4.(a), a set of random boxes and multi-scaled feature maps is first taken from the encoder as input. Then, dynamic queries (Figure 4b) are initialized from a regional pooling operator (RoIAlign [26]) and a feed-forward network (FFN) [57]. These queries undergo iterative refinement through multiple stages. Each stage takes proposal boxes and dynamic queries from the previous stage, generating refined dynamic queries, box predictions, and class predictions for the next stage. Figure 4c illustrates one detailed refinement stage. First, new regional features (cropped by RoIAlign) interact with dynamic queries (after a self-attention module) via dynamic convolution [11], emphasizing regions likely containing objects while suppressing others. Next, the enhanced regional features are processed by FFNs to produce refined box predictions, box-wise classifications, and queries, respectively.
@@ 3.2.3 Sampling
Conventional RCNN models use intersection over union (IoU) to classify a proposal box as positive if its IoU with a ground-truth box is $\geq 0.5$ [26]. This approach, however, is ill-suited for diffusion methods, particularly in early training stages when diffusion methods predominantly produce noise. That can complicate convergence or even cause failure due to error accumulation. Further, conventional IoU-based methods are susceptible to class imbalances, biasing the learning process towards frequently occurring objects, thereby neglecting rare instances. We, therefore, propose a novel sampling method to address these issues. We first replace proposal boxes with ground-truth boxes, ensuring a sufficient supply of positive samples. That is justified because the bbox-decoder aims to predict boxes that closely approximate the ground-truth. We then divide ground-truth bounding boxes into groups based on their classes. Within each group, we randomly select up to $n$ boxes as positive samples without repetition. The core idea is that objects of each class having similar shapes and appearances can be mapped to a cluster. Consequently, it allows us to learn instance feature representations with a small sample number $n$. This especially applies to the tubular class that occupies most of our dataset. This sampling strategy offers three distinct advantages:
@@
    Randomize Frequent Instance: Learning on a randomized subset of frequent instances prevents model over-fitting.
@@ 3.2.4 Mask decoder
The mask decoder iteratively refines mask predictions, utilizing regional features cropped from multi-scale feature maps based on boxes selected by the sampling module. Instead of using static queries [12, 13], we compute cross-attention between dynamic queries and positive regional features. That highlights effective queries and filters others, enhancing regional features by focusing on pixels within positive regions, thereby facilitating accurate per-pixel instance score prediction (Figure 5).
Fig. 4: The bounding box (bbox)-decoder takes multi-scale feature maps and a set of random boxes as input. Then, the prediction of classes and boxes will be outputted iteratively. (a) The module comprises an initialization head for dynamic queries (orange rectangle) and multiple box refinement heads (blue rectangles). (b) In the orange rectangle in (a), the initial dynamic queries are generated via a RoIAlign pooling operator and a feed-forward network (FFN). (c) Each box refinement module (one blue rectangle in (a)) takes the previous stage’s dynamic queries and proposal boxes as input, generating predictions and refined dynamic queries for the next stage. Abbreviations: dynamic queries: DQ, feed-forward network: FFN, regional features: RF.
Fig. 5: The mask-decoder takes multi-scale feature maps and a set of positive boxes as input and predicts instance masks. The dynamic queries interact with regional features using cross-attention and only highlight pixels residing in proposal boxes. Final instance masks are generated from the enhanced regional feature maps. Abbreviations: dynamic queries: DQ, feed-forward network: FFN, regional features: RF.
Fig. 6: The lesion classifier consists of 2 prediction heads, one dedicated to glomeruli and the other to tubuli. It takes croppings of anatomical structures and outputs the probability of class-wise lesions.
Fig. 7: An illustration for a diffusion model at different stages depicts a simplified process with $5$ states (skip step=$1$), where state $4$ is pure Gaussian noise and state 00 is the ground-truth. (a) During training, a diffusion model randomly selects denoising paths to train its decoder; (b) In inference, the Denoising Diffusion Implicit Model (DDIM) flexibly alternates between denoising and corrupting paths, generating multiple predictions that are then ensembled for a refined final result.
@@ 3.3 Lesion classifier
The lesion classifier comprises multiple independent lesion prediction heads. Each head specializes in a specific anatomical structure (e.g., glomerulus) and processes cropped images of the specified structure to predict probabilities for multi-label lesions. Figure 6 illustrates this design. The architecture of a single head consists of three convolutional blocks followed by a max-pooling operator. Each block contains a $3\times 3$ convolutional, batch normalization, and ReLU activation. We currently implement two prediction heads for glomeruli (predicting sclerosis) and tubuli (predicting atrophy).
Notably, the modular design of the instance segmentation component allows lesion prediction heads to operate as independent plug-ins. This modularity is crucial for seamlessly adapting the model to the expanding datasets, particularly in clinical scenarios. It allows for the straightforward integration or removal of prediction heads with minimal impact on the overall model architecture. For example, if new annotations for arteries become available, a dedicated prediction head can be trained to classify arterial lesions and integrate them into the model without modifying existing components. Likewise, this design extends the model’s diagnostic scope to include additional lesions within a given structure (e.g., tubuli) by simply replacing an existing prediction head with a newly trained one on expanded data. This design promotes the model’s versatility and long-term adaptability within clinical scenarios where the scope of identifiable lesions may increase.
@@ 3.4 Implementation Details
The entire model adopts distinct strategies for the training and inference phases. During training, the DiffRegFormer learns to reconstruct ground-truth boxes through adaptation at state 00 from an arbitrary noisy state $t$ (Figure 7.(a)). Additionally, the sampling module provides class-balanced ground-truth boxes to facilitate the mask decoder’s training. Lastly, The lesion classifier operates on croppings generated from ground-truth instance masks. Each prediction head specializes in classifying lesions specific to the corresponding anatomical structure within these croppings. At the same time, during inference, the DiffRegFormer assembles multiple predictions to refine final results (Figure 7b). Moreover, the sampling module is deactivated, and all proposals are directly forwarded to the mask decoder for instance segmentation prediction. Likewise, the lesion classifier utilizes instance masks predicted by DiffRegFormer to crop structures. As a last step, each prediction head predicts probabilities for class-specific lesions within these cropped regions.
def train(images, gt_boxes, gt_masks):
␣␣images:[B,H,W,3]
␣␣gt_boxes:[B,*,4]
␣␣gt_masks:[B,H,W,G]
␣␣B:batchsize
␣␣N:numberofproposalboxes
# generate multi-scale features via encoder
feats = encoder(images)
# generate noised boxes at state t where t
# is a random integer in diffusion forward;
# Pad boxes to N with t_boxes: [B, N, 4]
t_boxes, t = diffusion(gt_boxes, mean=0, std=1)
# generate initial dynamic queries
d_query = intialize_query(t_boxes, feats)
# learn to reverse noised boxes at state t back
# to ground-truth boxes at state 0 and return
# refined dynamic queries
[pred_boxes, d_query] = box_decoder(t_boxes, feats, d_query, t)
# obtain bbox loss via the set objective function
loss_bbox = set_loss(pred_boxes, gt_boxes)
# randomly select balanced positive boxes from
# groud-truth boxes
pos_boxes = sampling(gt_boxes)
# generate predicted instance masks
pred_masks = mask_decoder(pos_boxes, feats, d_query)
# obtain mask loss via the set objective function
loss_mask = set_prediction_loss(pred_masks, gt_masks)
return loss_bbox, loss_mask
Algorithm 1 DiffRegFormer Training
@@ 3.4.1 Training
We start by constructing noisy boxes from the ground-truth. The DiffRegFormer is then trained to recover bounding boxes and iteratively generate instance masks from an arbitrary state $t$. The algorithm is detailed in Algorithm 1.
Diffusion forward. Ground-truth boxes are first concatenated with additional random boxes to align to a fixed length across images, accounting for the variability in instance counts. Gaussian noise is then introduced as the noise scale governed by $\bar{\alpha}$ as per Eq.(1). Following [45], the scale varies with state $t$ using a monotonically decreasing cosine schedule. It should be noted that, for optimal performance, the signal-to-noise ratio requires a relatively high signal scaling value compared to image generation tasks [9, 29, 18, 10, 8].
Training Loss. The box decoder inputs noisy boxes and predicts category classifications and box coordinates. Given the one-to-many relationship between ground-truth and predicted boxes, first, a set-loss function is employed to match predictions to the ground-truth based on the lowest cost [5, 55, 62]. Furthermore, the mask decoder inputs sampled positive boxes and predicts instance masks. We use a Binary Cross Entropy loss function due to the one-to-one mapping. Finally, since multiple lesions can exist within one cropping, lesion classification adopts a Binary Cross Entropy loss function for multi-label prediction within each head.
Sampling. Our implementation uses $3$ classes and a maximum of $N=500$ objects per image. We set up to $n=\frac{500}{3}\simeq 166$ instances per class. More specifically, positive ground-truth boxes are sampled without repetition, not exceeding the class-wise maximum. Moreover, we disregard absent classes within the input. For example, if there are 2 glomeruli, 0 vessels, and 200 tubuli ground-truth boxes within an ROI, then the sampling module would select 2 glomeruli, 0 vessels, and 166 random tubuli as positive samples without duplication.
def infer(images, step, T):
␣␣images:[B,H,W,3]
␣␣step:theskiplengthforthestatetransform
␣␣T:thelengthofthechain
␣␣B:batchsize
␣␣N:numberofproposalboxes
# generate multi-scale features via encoder
feats = encoder(images)
# return Gaussian noise as noisy boxes at state T;
# Pad boxes to N with T_boxes: [B, N, 4]
[t_boxes, _] = diffusion(mean=0, std=1)
# generate state transform pairs skipping every step
# [(T, T-step), (T-step, T-2*step), ..., (step, 0)]
time_pairs = uniform(0, T, step)
# generate initial dynamic queries
d_query = intialize_query(t_boxes, feats)
# iterate over stages
for (t_now, t_next) iterate t_paris:
# predict boxes and dynamic query at state t_now
[pred_boxes, d_query] = box_decoder(t_boxes, feats, d_query, t_now)
# generate new noisy boxes from state t_now to t_next
t_boxes = ddim(t_boxes, pred_boxes, t_now, t_next)
# replace undesired boxes with random Gaussian noise
t_boxes = box_repalce(t_boxes)
# generate predicted instance masks
pred_masks = mask_decoder(pred_boxes, feats, d_query)
return pred_boxes, pred_masks
Algorithm 2 DiffRegFormer Inference
@@ 3.4.2 Inference
During inference, the DiffRegFormer starts from a standard Gaussian distribution corresponding to the final state $T$ as defined in Eq.(1). Then, a progressive denoising operation reverses the predictions to the initial state 00. The algorithm, as depicted in Algorithm 2, outlines this process.
Inference steps. Each inference iteration involves two primary operations. First, the box decoder processes noisy boxes generated from the previous state $\mathrm{t_{now}}$ and predicts both categories and box coordinates. Subsequently, a Denoising Diffusion Implicit Model (DDIM) [44, 53] introduces noise into the previously predicted boxes. This step generates new noisy boxes for subsequent state $\mathrm{t_{next}}$, facilitating a progressive denoising process across different states. Notably, each iteration performs a single denoising operation, transitioning from state $\mathrm{t_{now}}$ to $\mathrm{t_{next}}$. This characteristic allows for the assembly of intermediate predictions from multiple denoising steps (i.e., $t_{i}\to 0$ where $i\in[1,T]$) to refine the outcome. However, DiffRegFormer performs inference only once to improve efficiency, directly transitioning from pure Gaussian noise at state $T$ to state 00 to obtain the final predictions.
Box replacement. Predicted boxes can be categorized as either positive (scores above a threshold, containing objects of interest) or negative (scores below the threshold, arbitrarily located). Directly applying negative boxes into DDIM would significantly degrade the quality of newly generated noisy boxes. This issue stems from the fact that negative boxes, originating from corrupted boxes during the training phase, significantly deviate from the Gaussian distribution. To ensure consistency between the training and inference phases, we substitute negative boxes with random boxes sampled from a Gaussian distribution.
Table 1: Overall evaluation results. All models use ResNet-50 as a backbone network and have been trained 40000 iterations on an NVIDIA GeForce RTX 3090 GPU with a batch size of 2. APS, APM, and APL correspond to AP on small, medium, and large-sized objects, respectively. The best results are highlighted in bold. CMask-RCNN represents Cascade Mask-RCNN.
@@
Table 2: Per-class evaluation results. All models use ResNet-50 as a backbone network and have been trained 40000 iterations on one NVIDIA GeForce RTX 3090 GPU with a batch size of 2. APS, APM, and APL correspond to mAP on small, medium, and large-sized objects, respectively. The best results are highlighted in bold. CMask-RCNN represents Cascade Mask-RCNN.
Bounding Boxes - Glomeruli Instances - Glomeruli
@@
CMask-RCNN 23.0 38.0 26.9 - 17.8 33.3 20.1 39.4 21.9 - 16.9 27.5
our model DiffRegFormer 25.7 46.9 25.8 13.5 23.9 31.6 23.4 48.8 20.2 20.2 20.4 30.1